OpenStudy (anonymous):

1.) Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14 (0, 0) (6, 14) Infinitely Many Solutions No Solutions 2.) What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 -7 7 9 -9 3.) Which point lies in the solution set for the following system of inequalities? y < 2x + 4 y < -2x + 2 (1, 0) (-5, -2) (0, -3) (-1, 5) Please try to help. Any help from anyone is appreciated...

OpenStudy (anonymous):

1. no solutions. (3x-2y=6)-2 6x-4y=14 -6x+4y=12 6x-4y=14 then subtract or add down. 0=24. not equal, so no solution

OpenStudy (anonymous):

2. 9 5x+4y=1 4x+3y=-1 need to get rid of x to have y left (5x+4y=1)4 (4x+3y+-1)-5 20x+16y=4 -20x-15y=5 1y=9 y=9

OpenStudy (anonymous):

3. just plug in each set of coordinates at a time and see which one works